OpenStudy (anonymous):

please help!!!! ---differential equations: x'=3x-2y and y'=2x-2y

OpenStudy (anonymous):

Ok, and then what?

OpenStudy (anonymous):

solving simultaneously

OpenStudy (anonymous):

For x and y? Is x' = dx/dy and y' = dy/dx?

OpenStudy (anonymous):

yes... that's correct

OpenStudy (anonymous):

Hmm, I can see how to get an equation in x and y from that, but do you want numerical values for both x and y? I'll have to look at this a different way. Just a sec.

OpenStudy (anonymous):

The first thing I thought of was that x' and y' are reciprocals, so you could do \(\large 3x-2y=(2x-2y)^{-1}\) and rearrange from there...

OpenStudy (anonymous):

still confused...

OpenStudy (anonymous):

Me too. I'm not sure exactly what you're looking for.

OpenStudy (anonymous):

Is there any other information or detail in the question for what the answer should look like?

OpenStudy (anonymous):

an example in class, she said solve one eqn... let's say y=.. differentiate it and substitute in eqn (2)

OpenStudy (anonymous):

Ok, that sounds reasonable, being standard system solving procedure. That would look like this: \(\large x'=3x-2y \rightarrow y=(-1/2)(x'-3x)\) \(\large y'=2x-2y \rightarrow x=(1/2)(y'+2y)\) Differentiating x: \(\large x'=(1/2)(y''+2)\) Sub into other equation: \(\large y=(-1/2)((1/2)(y''+2)-3(1/2)(y'+2y))\) Simplify and rearange into a second-order DE. \(\large \)